# Canonical decompositions of 3-connected graphs

### v2

@@ 2.6. Proof of (iii)
Let $G$ be a graph, and let $\sigma$ be a star of ${}^{+}$ of $G$. Then every edge of $G$ lies in the separators of at most two elements of $\sigma$.
@@
This follows immediately from the observation that the vertex sets $A\setminus B$ are disjoint for distinct elements $(A,B)\in\sigma$. ∎
Figure 2.6. $\sigma$ and $\dot{\sigma}$
@@
Suppose now that $G$ is a graph and $\sigma$ is a star of ${}^{+}$ of $G$. In this context, we define the graph $\dot{G}$ to be the graph obtained from $G$ by subdividing every edge that lies in the separators of exactly two elements of $\sigma$. Let $(A,B)\in\sigma$ be arbitrary. We define $\dot{A}$ to be the vertex set obtained from $A$ by adding for every edge $e\in S(A,B)$ the subdividing vertex of $e$ if existent and the endvertex of $e$ in $B$ otherwise. We define $\dot{B}$ to be the vertex set obtained from $B$ by adding for every edge $e\in S(A,B)$ its subdividing vertex if existent (the endvertex of $e$ in $A$ is not added). Then $(\dot{A},\dot{B})$ is a ${}^{+}$ of $\dot{G}$, which has the same order as $(A,B)$. We write $\dot{\sigma}:=\{\,(\dot{A},\dot{B}):(A,B)\in\sigma\,\}$.
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Let $X$ denote the compressed torso of $\sigma$. We define a map $\iota\colon V(X)\to V(\dot{G})$ as follows. Let $v$ be any vertex of $X$. If $v$ is not a contraction vertex, then $v$ also is a vertex of $G$ and we let $\iota(v):=v$. Otherwise $v$ is a contraction vertex with branch set $U$. If $U$ is spanned in $G$ by a single edge $e$ that lies in the separators of two elements of $\sigma$, then we let $\iota$ take $v$ to the subdividing vertex of $e$ in $\dot{G}$. Else $U$ intersects the bag of $\sigma$ in a unique vertex, by Section 1.2, and we let $\iota$ take $v$ to this unique vertex. Let $\dot{\iota}$ be the restriction of $\iota$ onto its image.
@@
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Then $\dot{\sigma}$ is a star of ${}^{+}$ of $\dot{G}$, and $\dot{\iota}$ is a graph-isomorphism between the compressed torso of $\sigma$ in $G$ and the torso of $\dot{\sigma}$ in $\dot{G}$.∎
@@
Let $\sigma$ be a star of ${}^{+}$ of a graph $G$ and let $X$ denote the torso of $\sigma$. For every separation $(A,B)$ of $X$ there exists a separation $(\hat{A},\hat{B})$ of $G$ such that $\hat{A}\cap V(X)=A$ and $\hat{B}\cap V(X)=B$ and $\hat{A}\cap\hat{B}=A\cap B$. Moreover, $(\hat{A},\hat{B})$ almost interlaces $\sigma$.
@@
Let $(A,B)$ be given. For every $(C,D)\in\sigma$, the separator $C\cap D\subseteq V(X)$ is complete in $X$, so $C\cap D$ is included in $A$ or in $B$ (possibly in both). We obtain $\hat{A}$ from $A$ by adding all vertices in $C\setminus D$ from elements $(C,D)\in\sigma$ with $C\cap D\subseteq A$, and we obtain $\hat{B}$ from $B$ by adding all vertices in $C\setminus D$ from elements $(C,D)\in\sigma$ with $C\cap D\not\subseteq A$. Then $\hat{A}\cup\hat{B}=V(G)$. Let us assume for a contradiction that $G$ contains an edge $ab$ with $a\in\hat{A}\setminus\hat{B}$ and $b\in\hat{B}\setminus\hat{A}$. Since $(A,B)$ is a separation of $X$, not both $a$ and $b$ can lie in $V(X)$. So $a\in C\setminus D$ for some $(C,D)\in\sigma$ with $C\cap D\subseteq A$, say. Since $(C,D)$ is a ${}^{+}$, it follows that $b$ must lie in $C$, contradicting that $C\subseteq\hat{A}$.
The equalities $\hat{A}\cap V(X)=A$ and $\hat{B}\cap V(X)=B$ are immediate from the fact that $C\setminus D$ avoids $V(X)$ for all $(C,D)\in\sigma$. The equality $\hat{A}\cap\hat{B}=A\cap B$ follows from the fact that $C\setminus D$ is disjoint from $C^{\prime}\setminus D^{\prime}$ for every distinct two $(C,D),(C^{\prime},D^{\prime})\in\sigma$. ∎
In the context of Section 2.6, we say that $(A,B)$ lifts to $(\hat{A},\hat{B})$, and call $(\hat{A},\hat{B})$ a lift of $(A,B)$.
@@
If $G$ is a subdivision of a 3-connected graph, and $\sigma$ is a star of ${}^{+}$ of $G$, and the bag of $\sigma$ does not include a degree-two vertex plus both its neighbours, then the torso of $\sigma$ is 3-connected or a $K_{3}$.∎
@@
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Let $(C,D)$ be a separation of the compressed torso $X$ of $\sigma$. A hyper-lift of $(C,D)$ to $G$ is a mixed-separation $(\hat{C},\hat{D})$ of $G$ that is obtained from $(C,D)$ in the following way. First, we view $(C,D)$ as a separation of the torso of $\dot{\sigma}$, using Section 2.6. Next, we lift $(C,D)$ from the torso of $\dot{\sigma}$ to a separation $(C^{\prime},D^{\prime})$ of $\dot{G}$, using the Proof of (iii) (2.6). Finally, we let $\hat{C}:=C^{\prime}\cap V(G)$ and $\hat{D}:=D^{\prime}\cap V(G)$.
@@
(1). First, we shall define an injection from $S(\hat{C},\hat{D})$ to $S(C^{\prime},D^{\prime})$. To this end, let $x$ be an edge in $S(\hat{C},\hat{D})$. Since $(C^{\prime},D^{\prime})$ has no edges in its separator, the edge $x$ must be an edge of $G$ that is not an edge of $\dot{G}$. Denote by $y$ the unique subdivision vertex of $x$ in $\dot{G}$. Since the two neighbours of $y$ are in each of $C^{\prime}\setminus D^{\prime}$ and $D^{\prime}\setminus C^{\prime}$, and $S(C^{\prime},D^{\prime})$ contains no edges, the vertex $y$ is in $S(C^{\prime},D^{\prime})$. Let $\varphi(x):=y$. We extend $\varphi$ to a map from $S(\hat{C},\hat{D})$ to $S(C^{\prime},D^{\prime})$ by taking the identity on vertices. This map is injective since subdivision vertices $y\in V(\dot{G})$ of distinct edges of $G$ are distinct. Hence the order of $(\hat{C},\hat{D})$ is at most the order of $(C^{\prime},D^{\prime})$. By Section 2.6, the order of $(C^{\prime},D^{\prime})$ is equal to the order of $(C,D)$.
(2). We prove the stronger statement that $(C^{\prime},D^{\prime})$ almost interlaces $\dot{\sigma}$ (which gives the desired result as $V(G)\subseteq V(\dot{G})$ and we just need to restrict sides). This follows from Section 2.6.
@@
(4). Let $\beta$ denote the bag of $\dot{\sigma}$ in $\dot{G}$, and let $(A,B)\in\sigma$. By Section 2.6, we have that $|(C^{\prime}\setminus D^{\prime})\cap\beta|=|C\setminus D|$. We define an injection from $(C^{\prime}\setminus D^{\prime})\cap\beta$ to $(\hat{C}\setminus\hat{D})\cap B$. Assume that $v\in C^{\prime}\setminus D^{\prime}$ is a subdivision vertex of an edge $e$ of $G$ (so in particular $v\in\beta$). Then there is some $(E,F)\in\sigma$ that is different from $(A,B)$ that has the edge $e$ in its separator. Let $x$ be the endvertex of $e$ in $E\setminus F$.
Claim 2.6.0.1.
@@
Let $\varphi$ denote the map from $(C^{\prime}\setminus D^{\prime})\cap\beta$ to $(\hat{C}\setminus\hat{D})\cap B$ that is the identity on vertices of $G$ and maps subdivision vertices $v$ to vertices $x$ as defined above. Since $x$ is not in $\beta$, the sets $E^{\prime}\setminus F^{\prime}$ for $(E^{\prime},F^{\prime})\in\sigma$ are disjoint, and $x$ is incident with at most one edge of $S(E,F)$ by Section 1.2, the map $\varphi$ is injective. ∎
@@
By Section 2.4, we have $(A,B)\leqslant(C^{\prime},D^{\prime})$. By assumption, there is a vertex $v\in C\setminus D$ that lies in $B$. Then $v$ also lies in $(C^{\prime}\setminus D^{\prime})\cap B$. Hence the inclusion $B\supseteq D^{\prime}$ is proper. ∎
@@
Let $(\hat{C},\hat{D})$ be a hyper-lift of $(C,D)$ to $G$. By Section 2.6 (applied to $(C,D)$ and $(D,C)$), $(\hat{C},\hat{D})$ is a mixed 3-separation of $G$ that almost interlaces $\sigma$, and it satisfies $|\hat{C}\setminus\hat{D}|\geqslant 2$ and $|\hat{D}\setminus\hat{C}|\geqslant 2$ by (3). And by (4), for every $(A,B)\in\sigma$ we have $|(\hat{C}\setminus\hat{D})\cap B|\geqslant 2$ and $|(\hat{D}\setminus\hat{C})\cap B|\geqslant 2$. Let $(\bar{C},\bar{D})$ be a strengthening of $(\hat{C},\hat{D})$. Since $\bar{C}\setminus\bar{D}=\hat{C}\setminus\hat{D}$ and $\bar{D}\setminus\bar{C}=\hat{D}\setminus\hat{C}$, it follows with Section 1.2 that $(\bar{C},\bar{D})$ is nontrivial. By Section 2.4, $(\bar{C},\bar{D})$ almost interlaces $\sigma$. By Section 2.6, neither $(\bar{C},\bar{D})$ nor $(\bar{D},\bar{C})$ lies in $\sigma$. Hence $(\bar{C},\bar{D})$ interlaces $\sigma$. ∎
Proof of Theorem 2.2.1 (iii).
Let $G$ be a 3-connected graph, let $N$ denote its set of totally-nested nontrivial tri-separations, and let $\sigma$ be a splitting star of $N$. Suppose that $\sigma$ is not interlaced by a strong nontrivial tri-separation of $G$. We denote by $X$ the compressed torso of $\sigma$. By Section 2.6 and Section 2.6, $X$ is 3-connected or a $K_{3}$, and we are done in the latter case. The tri-separations in $\sigma$ are strong by Section 1.3. Hence by the contrapositive of Section 2.6, every 3-separation of $X$ has a side with at most four vertices. Hence $X$ is quasi 4-connected or a $K_{4}$. ∎
Proof of Theorem 2.2.1.
We have proved (i), (ii) and (iii) in the respective sections above. The ‘Moreover’ part holds by Section 2.2.3. ∎
@@
Theorem 2.2.1 implies Theorem 1. ∎
2.7. Tutte’s Wheel Theorem
@@
Theorem 2.7.1 (Tutte’s Wheel Theorem [62]).
@@
In this section we give an automatic proof of Tutte’s wheel theorem; the proof strategy is as follows. Take a minimally 3-connected graph $G$. First, we show that all totally-nested tri-separators of $G$ consist of three vertices that do not span any edge. Now consider a ‘leaf-torso’ [^2] A leaf-torso means a compressed torso of a splitting star $\{\,(A,B)\,\}$ where $(A,B)$ is a $\leqslant$-maximal totally-nested nontrivial tri-separation. of the set of totally-nested nontrivial tri-separations. By Theorem 2.2.1 there are three options how this torso might look like and one easily checks that none of them is possible. Hence $G$ has no totally-nested nontrivial tri-separation, and again by Theorem 2.2.1 we have three options how $G$ might look like; two are excluded for the same reasons and thus the only possibility is that $G$ is a wheel. The details are as follows.
@@
    if the ends of $e$ do not lie in the same 3-separator of $G$ and $e$ does not lie in a triangle of $G$, then $G/e$ is 3-connected;
@@
An edge $e$ of $G$ is of type c or d if it satisfies the premises of conditions (c) or (d) in Section 2.7, respectively.
@@
Let $G$ be a minimally 3-connected graph. The separator of every totally-nested tri-separation $(C,D)$ of $G$ consists of three vertices that do not span any edge.
@@
Let $U$ denote the set of vertices of $S(C,D)$ together with the endvertices of edges in $S(C,D)$. We claim that every edge $e=vw\notin S(C,D)$ between two vertices of $U$ is of type d. Indeed, the reduction $(E,F)$ of any mixed 3-separation with $e$ in the separator crosses $(C,D)$ with $v$ and $w$ in opposite links, which is not possible by total-nestedness of $(C,D)$. As $G$ is minimally 3-connected, the claim follows by Section 2.7.
We claim that every edge $e=vw$ in $S(C,D)$ is of type c. Indeed, by the above $e$ does not lie in a triangle. Moreover, the reduction $(E,F)$ of any mixed 3-separation with $v$ and $w$ in the separator also contains $v$ and $w$ in its separator (since $vw$ is an edge), and hence $(E,F)$ crosses $(C,D)$ with $v$ and $w$ in opposite links, which is not possible by total-nestedness of $(C,D)$. As $G$ is minimally 3-connected, $S(C,D)$ does not contain any edge by Section 2.7. This completes the proof. ∎
@@
For this, let $e$ be an arbitrary edge of $G$ with an endvertex in $X$. Since $e$ is not of type d by Section 2.7, it lies in the separator of a mixed 3-separation $(A,B)$ of $G$. Since we are done otherwise, we may assume that the reduction of $(A,B)$ is trivial. So an endvertex $v$ of $e$ has degree three. The lemma is trivial for $G=K_{4}$. So assume that there is a mixed 3-separation $(C,D)$ of $G$ with separator equal to $N(v)$.
Claim 2.7.1.1.
@@
Since we are done otherwise by Claim 2.7.1.1, assume that $N(v)$ does not span an edge; that is, $e$ is not in a triangle. Since $e=vw$ is not of type c by Section 2.7, there is a 3-separation $(E,F)$ with $v$ and $w$ in its separator; its reduction has the neighbours $v$ and $w$ in its separator and hence is nontrivial by Section 1.2. As one of $v$ and $w$ is in $X$, this gives the desired result. ∎
@@
The nontrivial tri-separation $(U,W)$ in Section 2.7 can be chosen strong.
@@
Via Section 1.2, take a strengthening of the tri-separation given by Section 2.7. ∎

### v3

@@ 6 Proof of (iii)
Let $G$ be a graph, and let $\sigma$ be a star of mixed-separations+ of $G$. Then every edge of $G$ lies in the separators of at most two elements of $\sigma$.
@@
Assume for a contradiction that some edge $e$ of $G$ lies in the separators of three elements $(A_{i},B_{i})$ of $\sigma$ for $i\in[3]$. Let $x\in A_{1}\setminus B_{1}$ and $y\in B_{1}\setminus A_{1}$ denote the endvertices of $e$. Since $(A_{1},B_{1})\leqslant(B_{j},A_{j})$ for both $j=2,3$ by the definition of star, the endvertex of $e$ in $A_{j}\setminus B_{j}$ must be $y$. So $A_{2}\setminus B_{2}$ and $A_{3}\setminus B_{3}$ intersect in the vertex $y$. But $(A_{2},B_{2})\leqslant(B_{3},A_{3})$ by the definition of star, which implies that $A_{2}\setminus B_{2}$ and $A_{3}\setminus B_{3}$ are disjoint, a contradiction. ∎
Figure 19: $\sigma$ and $\dot{\sigma}$
@@
Suppose now that $G$ is a graph and $\sigma$ is a star of mixed-separations+ of $G$. In this context, we define the graph $\dot{G}$ to be the graph obtained from $G$ by subdividing every edge that lies in the separators of exactly two elements of $\sigma$. Let $(A,B)\in\sigma$ be arbitrary. We define $\dot{A}$ to be the vertex set obtained from $A$ by adding for every edge $e\in S(A,B)$ the subdividing vertex of $e$ if existent and the endvertex of $e$ in $B$ otherwise. We define $\dot{B}$ to be the vertex set obtained from $B$ by adding for every edge $e\in S(A,B)$ its subdividing vertex if existent (the endvertex of $e$ in $A$ is not added). Then $(\dot{A},\dot{B})$ is a separation+ of $\dot{G}$, which has the same order as $(A,B)$. We write $\dot{\sigma}:=\{\,(\dot{A},\dot{B}):(A,B)\in\sigma\,\}$.
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Let $X$ denote the compressed torso of $\sigma$. We define a map $\iota\colon V(X)\to V(\dot{G})$ as follows. Let $v$ be any vertex of $X$. If $v$ is not a contraction vertex, then $v$ also is a vertex of $G$ and we let $\iota(v):=v$. Otherwise $v$ is a contraction vertex with branch set $U$. If $U$ is spanned in $G$ by a single edge $e$ that lies in the separators of two elements of $\sigma$, then we let $\iota$ take $v$ to the subdividing vertex of $e$ in $\dot{G}$. Else $U$ intersects the bag of $\sigma$ in a unique vertex, by Section 3, and we let $\iota$ take $v$ to this unique vertex. Let $\dot{\iota}$ be the restriction of $\iota$ onto its image.
@@
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Then $\dot{\sigma}$ is a star of 3-separations+ of $\dot{G}$, and $\dot{\iota}$ is a graph-isomorphism between the compressed torso of $\sigma$ in $G$ and the torso of $\dot{\sigma}$ in $\dot{G}$.∎
@@
Let $\sigma$ be a star of separations+ of a graph $G$ and let $X$ denote the torso of $\sigma$. For every separation $(A,B)$ of $X$ there exists a separation $(\hat{A},\hat{B})$ of $G$ such that $\hat{A}\cap V(X)=A$ and $\hat{B}\cap V(X)=B$ and $\hat{A}\cap\hat{B}=A\cap B$. Moreover, $(\hat{A},\hat{B})$ almost interlaces $\sigma$.
@@
Let $(A,B)$ be given. For every $(C,D)\in\sigma$, the separator $C\cap D\subseteq V(X)$ is complete in $X$, so $C\cap D$ is included in $A$ or in $B$ (possibly in both). We obtain $\hat{A}$ from $A$ by adding all vertices in $C\setminus D$ from elements $(C,D)\in\sigma$ with $C\cap D\subseteq A$, and we obtain $\hat{B}$ from $B$ by adding all vertices in $C\setminus D$ from elements $(C,D)\in\sigma$ with $C\cap D\not\subseteq A$. Then $\hat{A}\cup\hat{B}=V(G)$. Let us assume for a contradiction that $G$ contains an edge $ab$ with $a\in\hat{A}\setminus\hat{B}$ and $b\in\hat{B}\setminus\hat{A}$. Since $(A,B)$ is a separation of $X$, not both $a$ and $b$ can lie in $V(X)$. So $a\in C\setminus D$ for some $(C,D)\in\sigma$ with $C\cap D\subseteq A$, say. Since $(C,D)$ is a separation+, it follows that $b$ must lie in $C$, contradicting that $C\subseteq\hat{A}$.
The equalities $\hat{A}\cap V(X)=A$ and $\hat{B}\cap V(X)=B$ are immediate from the fact that $C\setminus D$ avoids $V(X)$ for all $(C,D)\in\sigma$. The equality $\hat{A}\cap\hat{B}=A\cap B$ follows from the fact that $C\setminus D$ is disjoint from $C^{\prime}\setminus D^{\prime}$ for every distinct two $(C,D),(C^{\prime},D^{\prime})\in\sigma$.
To see that $(\hat{A},\hat{B})$ almost interlaces $\sigma$, let $(C,D)\in\sigma$ be arbitrary. Since the separator $C\cap D$ is complete in $X$, we have $C\cap D\subseteq A$ or $C\cap D\subseteq B$. In the former case, we have $(C,D)\leqslant(\hat{A},\hat{B})$ since $C\subseteq\hat{A}$ by construction and $D\supseteq V(X)\cup\bigcup_{(C^{\prime},D^{\prime})\in\sigma\setminus\{(C,D)\}}D^{\prime}\supseteq\hat{B}$ by the definition of star and construction of $\hat{B}$. In the latter case, we similarly find that $(C,D)\leqslant(\hat{B},\hat{A})$. Therefore, $(\hat{A},\hat{B})$ almost interlaces $\sigma$. ∎
In the context of Section 6, we say that $(A,B)$ lifts to $(\hat{A},\hat{B})$, and call $(\hat{A},\hat{B})$ a lift of $(A,B)$.
@@
If $G$ is a subdivision of a 3-connected graph, and $\sigma$ is a star of 3-separations+ of $G$, and the bag of $\sigma$ does not include a degree-two vertex plus both its neighbours, then the torso of $\sigma$ is 3-connected or a $K_{3}$.∎
@@
Let $G$ be a 3-connected graph, and let $\sigma$ be a star of nontrivial tri-separations of $G$. Let $(C,D)$ be a separation of the compressed torso $X$ of $\sigma$. A hyper-lift of $(C,D)$ to $G$ is a mixed-separation $(\hat{C},\hat{D})$ of $G$ that is obtained from $(C,D)$ in the following way. First, we view $(C,D)$ as a separation of the torso of $\dot{\sigma}$, using Section 6. Next, we lift $(C,D)$ from the torso of $\dot{\sigma}$ to a separation $(C^{\prime},D^{\prime})$ of $\dot{G}$, using the Proof of (iii) (6). Finally, we let $\hat{C}:=C^{\prime}\cap V(G)$ and $\hat{D}:=D^{\prime}\cap V(G)$.
@@
(1). First, we shall define an injection from $S(\hat{C},\hat{D})$ to $S(C^{\prime},D^{\prime})$. To this end, let $x$ be an edge in $S(\hat{C},\hat{D})$. Since $(C^{\prime},D^{\prime})$ has no edges in its separator, the edge $x$ must be an edge of $G$ that is not an edge of $\dot{G}$. Denote by $y$ the unique subdivision vertex of $x$ in $\dot{G}$. Since the two neighbours of $y$ are in each of $C^{\prime}\setminus D^{\prime}$ and $D^{\prime}\setminus C^{\prime}$, and $S(C^{\prime},D^{\prime})$ contains no edges, the vertex $y$ is in $S(C^{\prime},D^{\prime})$. Let $\varphi(x):=y$. We extend $\varphi$ to a map from $S(\hat{C},\hat{D})$ to $S(C^{\prime},D^{\prime})$ by taking the identity on vertices. This map is injective since subdivision vertices $y\in V(\dot{G})$ of distinct edges of $G$ are distinct. Hence the order of $(\hat{C},\hat{D})$ is at most the order of $(C^{\prime},D^{\prime})$. By Section 6, the order of $(C^{\prime},D^{\prime})$ is equal to the order of $(C,D)$.
(2). We prove the stronger statement that $(C^{\prime},D^{\prime})$ almost interlaces $\dot{\sigma}$ (which gives the desired result as $V(G)\subseteq V(\dot{G})$ and we just need to restrict sides). This follows from Section 6.
@@
(4). Let $\beta$ denote the bag of $\dot{\sigma}$ in $\dot{G}$, and let $(A,B)\in\sigma$. By Section 6, we have that $|(C^{\prime}\setminus D^{\prime})\cap\beta|=|C\setminus D|$. We define an injection from $(C^{\prime}\setminus D^{\prime})\cap\beta$ to $(\hat{C}\setminus\hat{D})\cap B$. Assume that $v\in C^{\prime}\setminus D^{\prime}$ is a subdivision vertex of an edge $e$ of $G$ (so in particular $v\in\beta$). Then there is some $(E,F)\in\sigma$ that is different from $(A,B)$ that has the edge $e$ in its separator. Let $x$ be the endvertex of $e$ in $E\setminus F$.
Claim 6.0.1.
@@
Let $\varphi$ denote the map from $(C^{\prime}\setminus D^{\prime})\cap\beta$ to $(\hat{C}\setminus\hat{D})\cap B$ that is the identity on vertices of $G$ and maps subdivision vertices $v$ to vertices $x$ as defined above. Since $x$ is not in $\beta$, the sets $E^{\prime}\setminus F^{\prime}$ for $(E^{\prime},F^{\prime})\in\sigma$ are disjoint, and since $x$ is incident with at most one edge of $S(E,F)$ by Section 3, the map $\varphi$ is injective. ∎
@@
By Section 4, we have $(A,B)\leqslant(C^{\prime},D^{\prime})$. By assumption, there is a vertex $v\in C\setminus D$ that lies in $B$. Then $v$ also lies in $(C^{\prime}\setminus D^{\prime})\cap B$. Hence the inclusion $B\supseteq D^{\prime}$ is proper. ∎
@@
Let $(\hat{C},\hat{D})$ be a hyper-lift of $(C,D)$ to $G$. By Section 6 (applied to $(C,D)$ and $(D,C)$), $(\hat{C},\hat{D})$ is a mixed 3-separation of $G$ that almost interlaces $\sigma$, and it satisfies $|\hat{C}\setminus\hat{D}|\geqslant 2$ and $|\hat{D}\setminus\hat{C}|\geqslant 2$ by (3). Moreover, by (4), for every $(A,B)\in\sigma$ we have $|(\hat{C}\setminus\hat{D})\cap B|\geqslant 2$ and $|(\hat{D}\setminus\hat{C})\cap B|\geqslant 2$. Let $(\bar{C},\bar{D})$ be a strengthening of $(\hat{C},\hat{D})$. Since $\bar{C}\setminus\bar{D}=\hat{C}\setminus\hat{D}$ and $\bar{D}\setminus\bar{C}=\hat{D}\setminus\hat{C}$, it follows with Section 3 that $(\bar{C},\bar{D})$ is nontrivial. By Section 4, $(\bar{C},\bar{D})$ almost interlaces $\sigma$. By Section 6, neither $(\bar{C},\bar{D})$ nor $(\bar{D},\bar{C})$ lies in $\sigma$. Hence $(\bar{C},\bar{D})$ interlaces $\sigma$. ∎
Proof of Theorem 2.1 (iii).
Let $G$ be a 3-connected graph, let $N$ denote its set of totally-nested nontrivial tri-separations, and let $\sigma$ be a splitting star of $N$. Suppose that $\sigma$ is not interlaced by a strong nontrivial tri-separation of $G$. We denote by $X$ the compressed torso of $\sigma$. By Section 6 and Section 6, $X$ is 3-connected or a $K_{3}$, and we are done in the latter case. The tri-separations in $\sigma$ are strong by Section 4. Hence by the contrapositive of Section 6, every 3-separation of $X$ has a side with at most four vertices. Hence $X$ is quasi 4-connected or a $K_{4}$. ∎
Proof of Theorem 2.1.
We have proved (i), (ii) and (iii) in the respective sections above. The ‘Moreover’ part holds by Section 3. ∎
@@
Theorem 2.1 implies Theorem 1. ∎
7 Tutte’s Wheel Theorem
@@
Theorem 7.1 (Tutte’s Wheel Theorem [65]).
@@
In this section we give an automatic proof of Tutte’s wheel theorem; the proof strategy is as follows. Take a minimally 3-connected graph $G$. First, we show that all totally-nested tri-separators of $G$ consist of three vertices that do not span any edge. Now consider a ‘leaf-torso’ [^2] A leaf-torso means a compressed torso of a splitting star $\{\,(A,B)\,\}$ where $(A,B)$ is a $\leqslant$-maximal totally-nested nontrivial tri-separation. of the set of totally-nested nontrivial tri-separations. By Theorem 2.1 there are three options how this torso might look like and one easily checks that none of them is possible. Hence $G$ has no totally-nested nontrivial tri-separation, and again by Theorem 2.1 we have three options how $G$ might look like; two are excluded for the same reasons and thus the only possibility is that $G$ is a wheel. The details are as follows. We use ‘(c)’ and ‘(d)’ as a reminder for contraction and deletion.
@@
    if the ends of $e$ do not lie in a common 3-separator of $G$ and $e$ does not lie in a triangle of $G$, then $G/e$ is 3-connected;
@@
An edge $e$ of $G$ is of type c or d if it satisfies the premises of conditions (c) or (d) in Section 7, respectively.
@@
Let $G$ be a 3-connected graph. Let $(A,B)$ be a totally-nested nontrivial tri-separation of $G$. Assume that $S(A,B)$ contains an edge $ab$ with $a\in A\setminus B$ and $b\in B\setminus A$ such that $ab$ lies in a triangle $\Delta\subseteq G$. Let $A^{\prime}:=A\cup\{b\}$ and $B^{\prime}:=B$. Then $(A^{\prime},B^{\prime})$ is a totally-nested nontrivial tri-separation of $G$ with $(A,B)<(A^{\prime},B^{\prime})$.
@@
Let $x$ denote the third vertex in $\Delta$ besides $a$ and $b$, and note that $x\in S(A,B)$. The separator $S(A^{\prime},B^{\prime})$ is obtained from $S(A,B)$ by replacing the edge $ab$ with the vertex $b$; in particular, $x\in S(A^{\prime},B^{\prime})$. The edge $xb$ ensures that $(A^{\prime},B^{\prime})$ is a tri-separation, and clearly $(A^{\prime},B^{\prime})$ is nontrivial. Assume for a contradiction that $(A^{\prime},B^{\prime})$ is crossed by a tri-separation $(C,D)$. Since $(A,B)<(A^{\prime},B^{\prime})$, we must have $(A,B)\leqslant(C,D)$ or $(A,B)\leqslant(D,C)$; say $(A,B)\leqslant(C,D)$. The edge $ab$ witnesses that $G[A^{\prime}\setminus B^{\prime}]$ is connected, so $(A^{\prime},B^{\prime})$ is half-connected. Hence $(A^{\prime},B^{\prime})$ and $(C,D)$ cross with all links of size one and a single vertex $v$ in the centre, by the Nested or crossed: analysing corner diagrams (4).
We claim that $v=x$. Since $b,x\in S(A^{\prime},B^{\prime})$ and $bx$ is an edge in $G$, we have $v=b$ or $v=x$. We have $b\notin C$ since otherwise $A^{\prime}=A\cup\{b\}\subseteq C$ and $B^{\prime}=B\supseteq D$ (from $(A,B)\leqslant(C,D)$) together give $(A^{\prime},B^{\prime})\leqslant(C,D)$, a contradiction. Hence $ab\in S(C,D)$. So neither $a$ nor $b$ lies in $S(C,D)$. In particular, $v\neq b$, and therefore $v=x$.
Let $C^{\prime}:=C$ and $D^{\prime}:=D\cup\{a\}$. We claim that $(C^{\prime},D^{\prime})$ is a tri-separation that crosses $(A,B)$. The separator of $(C^{\prime},D^{\prime})$ is obtained from the separator of $(C,D)$ by replacing the edge $ab$ with the vertex $a$. The vertex $x\in S(C^{\prime},D^{\prime})$ is a neighbour of $a$, which ensures that $(C^{\prime},D^{\prime})$ is a tri-separation. To show that $(C^{\prime},D^{\prime})$ crosses $(A,B)$, it suffices to find two opposite links that are non-empty. The link for $A$ contains the vertex $a$. The link for $B$ contains the non-empty link for $B^{\prime}$ with regard to how $(C,D)$ crosses $(A^{\prime},B^{\prime})$, as $B=B^{\prime}$ and $\{a,b\}\subseteq A^{\prime}$. So $(C^{\prime},D^{\prime})$ crosses $(A,B)$, contradicting that $(A,B)$ is totally-nested. ∎
[…]
Let $G$ be a minimally 3-connected graph. Among the totally-nested nontrivial tri-separations of $G$, let $(C,D)$ be maximal with regard to the partial order $\leqslant$. Then $S(C,D)$ consists of three vertices that do not span any edge.
Proof.
We claim that every edge $e=vw$ in $S(C,D)$ is of type c. Indeed, by Section 7 and the maximality of $(C,D)$, the edge $e$ does not lie in a triangle. Moreover, the reduction $(E,F)$ of any mixed 3-separation with $v$ and $w$ in its separator also contains $v$ and $w$ in its separator (since $vw$ is an edge), and hence $(E,F)$ crosses $(C,D)$ with $v$ and $w$ in opposite links, which is not possible by total-nestedness of $(C,D)$. So every edge in $S(C,D)$ is of type c.
As $G$ is minimally 3-connected, $S(C,D)$ does not contain any edge by Section 7. So $S(C,D)$ consists of three vertices.
We claim that every edge $e=vw$ between two vertices $v,w\in S(C,D)$ is of type d. Indeed, the reduction $(E,F)$ of any mixed 3-separation with $e$ in its separator crosses $(C,D)$ with $v$ and $w$ in opposite links, which is not possible by total-nestedness of $(C,D)$.
As $G$ is minimally 3-connected, no two vertices in $S(C,D)$ span an edge by Section 7. This completes the proof. ∎
@@
For this, let $e$ be an arbitrary edge of $G$ with an endvertex in $X$. Since $e$ is not of type d by Section 7, it lies in the separator of a mixed 3-separation $(A,B)$ of $G$. Since we are done otherwise, we may assume that the reduction of $(A,B)$ is trivial. So an endvertex $v$ of $e$ has degree three. The lemma is trivial for $G=K_{4}$. So assume that there is a mixed 3-separation $(C,D)$ of $G$ with separator equal to $N(v)$.
Claim 7.1.1.
@@
Since we are done otherwise by Claim 7.1.1, assume that $N(v)$ does not span an edge; that is, $e$ is not in a triangle. Since $e=vw$ is not of type c by Section 7, there is a 3-separation $(E,F)$ with $v$ and $w$ in its separator; its reduction has the neighbours $v$ and $w$ in its separator and hence is nontrivial by Section 3. As one of $v$ and $w$ is in $X$, this gives the desired result. ∎
@@
The nontrivial tri-separation $(U,W)$ in Section 7 can be chosen strong.
@@
Via Section 3, take a strengthening of the tri-separation given by Section 7. ∎